### v2

# Market Efficiency and the Heterogeneous Impact of Financial Liberalization: Evidence from the Shanghai-Hong Kong Stock Connect

Abstract
This paper investigates the impact of the Shanghai-Hong Kong Stock Connect (SHHK Stock Connect) on the A-H share price premium and examines whether the policy effect is contingent on market efficiency. Using monthly data for 67 Shanghai-listed A-H dual-listed firms from January 2011 to May 2019, we employ a dynamic panel model estimated via two-step system generalized method of moment (GMM) to account for the persistence of the premium and potential endogeneity. Market efficiency is proxied by trading-friction measures derived from daily high-low price ranges. Our findings indicate that the implementation of SHHK Stock Connect is associated with an average 18.4% increase in the A-H premium. However, this effect is heterogeneous: the marginal impact of the policy is more pronounced for firms operating in less efficient markets and weaker for those with higher efficiency, suggesting that pre-existing trading frictions shape the policy outcome. No significant response is observed at the announcement stage. Placebo tests and alternative efficiency measures confirm the robustness of the efficiency-dependent effect. Overall, the results underscore the importance of the information environment in shaping the outcomes of financial liberalization.
Keywords: A-H premium; Stock Connect; market efficiency; capital market liberalization; cross-listed firms
JEL Classification: G14; G15; G18
1 Introduction
In 1993, Tsingtao Brewery became the first firm to achieve A-H dual listing, with shares traded on both the Shanghai Stock Exchange (SSE) and the Hong Kong (HKSE). [^2] An A-H dual-listed firm is a Chinese company with shares listed on both the mainland Chinese stock exchanges (A-shares, traded in CNY) and the Hong Kong Stock Exchange (H-shares, traded in HKD). This landmark listing marked the beginning of a growing trend of mainland Chinese companies pursuing dual listings across the two markets. Over the subsequent three decades, more than 100 firms have successfully completed A-H dual listings and raised capital through this channel.
According to the law of one price, the prices of dual-listed shares should be equal after adjusting for exchange rates; otherwise, arbitrage opportunities would exist. However, this theoretical implication does not hold in the case of Chinese A-H dual-listed firms, giving rise to what is commonly referred to as the A-H premium. Shares of the same company traded in the A-share and H-share markets frequently exhibit persistent price differentials. Early evidence from Bailey, (1994) documents that domestically listed shares are often priced at a premium relative to their internationally listed counterparts. Extending this analysis to the Chinese context, Chan and Kwok, (2005) examine A-shares, B-shares, and H-shares of Chinese firms and find that A-share prices are generally higher than both B-share and H-share prices. Similarly, Li et al., (2006) study A-H dual-listed firms and conclude that such companies predominantly exhibit a significant premium of A-shares over H-shares.
On 10 April 2014, the China Securities Regulatory Commission and (CSRC) the Securities and Futures Commission (SFC) jointly approved the pilot framework for the SHHK Stock Connect, establishing an institutional mechanism for cross-border equity trading between the Shanghai and Hong Kong markets. The program was jointly implemented by the Shanghai Stock Exchange, Hong Kong Exchanges and Clearing Limited, China Securities Depository and Clearing Corporation Limited, and Hong Kong Securities Clearing Company Limited, with trading formally commencing on 17 November 2014. As the first formal link between the equity markets of Mainland China and Hong Kong, the SHHK Stock Connect represented a pivotal step in China’s gradual capital-market liberalization. By enabling mutual market access for domestic and international investors within a regulated framework, the scheme deepened financial integration between the two markets and advanced the two-way opening of China’s capital account. In doing so, it reinforced the international roles of Shanghai and Hong Kong as major financial centers and contributed to the outward-oriented development of China’s capital markets. Moreover, by expanding cross-border investment channels, the SHHK Stock Connect has the potential to influence the A-H premium through changes in trading frictions and arbitrage capacity.
Following the launch of the SHHK Stock Connect, a growing body of research has examined its impact on the pricing of A-H dual-listed shares. Employing a vector error correction model (VECM), Chan and Kwok, (2016) analyze the A-H share price premium and find that while the announcement of the Stock Connect led to a significant reduction in the equilibrium premium, the premium widened sharply during the post-implementation period. Similarly, Hui and Chan, (2018) examine 12 representative dual-listed companies in Shanghai and Hong Kong and report a significant increase in the A-H premium following the implementation of the SHHK Connect. However, other studies offer contrasting evidence. Fan and Wang, (2017) introduce corporate governance as a mechanism to explain the A-H premium and, drawing on the market segmentation hypothesis, find that the opening of SHHK Stock Connect significantly narrows the price gap between A-shares and H-shares. In a related vein, Pan and Chi, (2021) employ the generalized method of moments (GMM) to analyze 56 Shanghai- and Hong Kong-listed dual firms and conclude that the SHHK Stock Connect significantly reduces the A-H share price premium.
Existing research has primarily focused on the average effect of the SHHK Stock Connect, yet the findings remain inconclusive. Building on the theoretical framework of market liberalization, a growing body of literature emphasizes that the impact of policy shocks is not uniform across all securities. Bekaert et al., (2005) find that the effects of financial liberalization depend on pre-existing levels of financial development and institutional quality. This heterogeneity is further supported by Mitton, (2006), who demonstrate that the positive impacts of stock market opening are significantly stronger for certain firms, depending on their specific characteristics and degree of ”investability.” Bond et al., (2012) argue that market prices can influence real outcomes through a feedback effect channel: when prices are more informative for real decision-making, economic agents that rely on market signals can respond more effectively, whereas limited informativeness weakens such transmission.
Taken together, the mixed findings in the literature suggest that the impact of the SHHK Stock Connect may vary across A-H dual-listed shares. Motivated by this possibility, our study examines not only the average effect of the SHHK Stock Connect on the A-H premium but also whether the policy impact varies systematically with the market efficiency of A- and H-share markets. This paper makes three contributions to the literature. First, we demonstrate that the effect of the SHHK Stock Connect on the A-H premium is heterogeneous, which helps reconcile the mixed findings in studies that focus solely on average effects. Second, following Meng et al., (2023) and Corwin and Schultz, (2012), we proxy market efficiency using a trading-frictions measure derived from high–low price ranges and find that the policy effects vary systematically with pre-existing market frictions. Third, we estimate the policy–premium relationship within a dynamic panel framework that accounts for the strong persistence of the A-H premium and corroborate the main findings through placebo tests based on alternative policy timing and robustness checks using alternative friction proxies.
The remainder of the paper is organized as follows. Section 2 details the data and methodology. Section 3 presents the empirical results. Section 4 concludes the paper, highlighting the limitation in the research.
2 Data and Methodology
2.1 Sample data
The sample period spans from January 2011 to May 2019, covering a total of 101 months. Following Allen et al., (2024), we select May 2019 as the endpoint to avoid potential confounding effects from the Anti-ELAB movement [^3] Anti-ELAB refers to the Anti-Extradition Law Amendment Bill Movement in Hong Kong. and the COVID-19 pandemic. We employ monthly rather than daily data for two primary reasons. First, daily observations are more susceptible to short-term fluctuations and transitory shocks, which can generate extreme values and introduce excessive noise into the analysis. Second, the trading calendars of the A-share and H-share markets differ, particularly due to disparities in public holidays between mainland China and Hong Kong. For instance, on a given calendar day, trading in the A-share market may correspond to the second trading day of the week, while trading in the H-share market may represent the first trading day following a holiday. Market sentiment can vary substantially under such conditions, and directly comparing daily prices across the two markets may therefore yield biased estimates. To mitigate this issue, we aggregate daily observations to the monthly level by averaging all daily observations within each month.
To identify the effect of the Shanghai–Hong Kong Stock Connect (SHHK), we restrict our sample to firms that completed their A–H dual listings prior to the implementation of the SHHK program and whose A-shares are listed on the Shanghai Stock Exchange. We also exclude firm-month observations with missing values. The final sample comprises 6,462 firm-month observations from 67 unique A-H dual-listed firms. Appendix A provides a complete list of sample firms, including company names, A-share listing dates and codes, and H-share listing dates and codes. A-share and H-share price data are obtained from the Wind Financial Terminal, while other share-related variables and firm-level characteristics are sourced from the Choice database.
2.2 Variable construction
Following Fan and Wang, (2017), we define the A-H share premium (Prem) as the ratio of the A-share price to the corresponding H-share price, adjusted for exchange rate differences. To account for potential autocorrelation in the premium series, we follow Fan and Wang, (2017) and Pan and Chi, (2021) by including the lagged premium term, $\texttt{Prem}_{t-1}$, in our dynamic panel specification. To capture the effect of the Shanghai–Hong Kong Stock Connect, we construct a binary indicator (SHHKPolicy) that equals one for the post-implementation period (i.e., after November 2014) and zero otherwise.
To evaluate the efficiency of the A-share and H-share markets, we employ a trading-frictions measure based on daily high and low prices. Specifically, we estimate the bid-ask spread using the method proposed by Corwin and Schultz, (2012), which derives spread estimates from the ratio of high-to-low prices over single and two-day intervals. Following the literature, we construct the following equations to compute the daily bid-ask spread.
$\mathbb{E}\left\{\sum_{j=0}^{1}\left[\ln\!\left(\frac{\text{H}^{0}_{t+j}}{\text{L}^{0}_{t+j}}\right)\right]^{2}\right\}=2k_{1}\sigma_{\text{HL}}^{2}+4k_{2}\sigma_{\text{HL}}\ln\!\left(\frac{2+\text{S}}{2-\text{S}}\right)+2\left[\ln\!\left(\frac{2+\text{S}}{2-\text{S}}\right)\right]^{2},$ (1)
and
$\mathbb{E}\left\{\left[\ln\!\left(\frac{\text{H}^{0}_{t,t+1}}{\text{L}^{0}_{t,t+1}}\right)\right]^{2}\right\}=2k_{1}\sigma_{\text{HL}}^{2}+2\sqrt{2}k_{2}\sigma_{\text{HL}}\ln\!\left(\frac{2+\text{S}}{2-\text{S}}\right)+\left[\ln\!\left(\frac{2+\text{S}}{2-\text{S}}\right)\right]^{2},$ (2)
where S denotes the bid-ask spread; $\text{H}^{0}_{t+j}$ and $\text{L}\textbf{}^{0}_{t+j}$ are the daily highest and lowest prices of a stock at the trading day of $t+j$ , respectively; $\text{H}^{0}_{t,t+1}$ and $\text{L}^{0}_{t,t+1}$ are the daily highest and lowest prices of a stock observed on two consecutive trading days $t$ and $t+1$, respectively; $\sigma_{HL}$ denotes volatility and $k_{1}$ and $k_{2}$ are coefficients.
It is not hard to see that Equation 1 is a one-day range of the spread (S) at the time of $t+j$. Equation 2 is a two-day range of the spread (S) from the time of $t$ to $t+1$.
Additionally, in order to make the process easier and in line with Corwin and Schultz, (2012), we define following:
$\beta=\mathbb{E}\left\{\sum_{j=0}^{1}\left[\ln\!\left(\frac{H^{0}_{t+j}}{L^{0}_{t+j}}\right)\right]^{2}\right\},\qquad\gamma=\mathbb{E}\left\{\left[\ln\!\left(\frac{H^{0}_{t,t+1}}{L^{0}_{t,t+1}}\right)\right]^{2}\right\}$ (3)
Solving Equations 1 and Equation 2, we can obtain the spread (S) equation, which is our closed form equation.
$S=\frac{2\left(e^{\alpha}-1\right)}{1+e^{\alpha}},$
where
$\alpha=\frac{\sqrt{2\beta}-\sqrt{\beta}}{3-2\sqrt{2}}-\sqrt{\frac{\gamma}{3-2\sqrt{2}}}.$
After computing daily bid-ask spreads, we aggregate them to the monthly level by taking the monthly average to align with the frequency of the other variables. [^4] For negative daily bid–ask spread estimates (S), we follow Corwin and Schultz, (2012) and set negative values to zero prior to calculating monthly averages. This approach yields more accurate estimates than either retaining negative spread observations or excluding them from the sample. In line with Meng et al., (2023), we use the negative of the bid-ask spread to proxy for market efficiency, implying an inverse relationship between the bid-ask spread and market efficiency. Additionally, we construct an aggregate efficiency measure, Effboth, by summing the A-share and H-share market efficiency proxies to capture the overall information environment of A-H dual-listed firms. [^5] It is worth noting that $Effboth$ is strictly negative in our sample by construction. Values closer to zero indicate higher market efficiency, while more negative values correspond to lower efficiency.
Regarding control variables, we account for differences in liquidity between the A-share and H-share markets by including the firm-level relative turnover ratio, Turnover, measured at monthly frequency, as a proxy for relative liquidity (Lee et al.,, 2001). Consistent with Chakravarty et al., (1998), firm size is proxied by market capitalization. To mitigate the influence of extreme observations arising from the positively skewed distribution of market capitalization, we take its natural logarithm. [^6] Market capitalization is calculated as the aggregate value of all shares listed across different stock exchanges and is adjusted for exchange rate differences. Firm size (Size) is included to capture differences in information asymmetry, as larger firms are generally associated with lower levels of asymmetric information.
We also include an indicator variable, SOE, which equals one if the firm is state-owned and zero otherwise. Following Chan and Kwok, (2005), we control for relative share demand (Demand) by incorporating the ratio of the number of floating A-shares to the number of floating H-shares.
Finally, to capture relative monetary conditions between the two markets, we construct a relative interest rate variable, Interest rate, defined as the ratio of the three-month Shanghai Interbank Offered Rate (SHIBOR) to the three-month Hong Kong Interbank Offered Rate (HIBOR).
Figure 1: Trend of AH premium from Jan 2011 to May 2019
2.3 Regression model
In this paper, we employ the generalized method of moments (GMM) rather than the traditional fixed-effects estimator. This approach has two main advantages. First, it allows us to control for serial correlation in the dependent variable by including its lagged value. Second, it helps mitigate potential endogeneity concerns by introducing internal instrumental variables. To test our hypothesis, we propose following regression framework:
$\text{Prem}_{i,t}=\alpha_{i}+\beta\,\text{Prem}_{i,t-1}+\theta\,\text{SHHKPolicy}_{t}+\gamma^{\prime}X_{it}+\varepsilon_{it}.$
where the $\text{Prem}_{i,t}$ is the A-H share price premium for share $i$ at month $t$, which is calculated as:
$\text{Prem}_{i,t}=\frac{P^{A}_{it}}{P^{H}_{it}\times\text{HKD/CNY}_{t}}-1$
Following Domowitz et al., (1997) and Fan and Wang, (2017), we control the lagged term $\text{Prem}_{i,t-1}$ to reduce the effect of autocorrelation. $\text{SHHKPolicy}_{t}$ is an indicator that equals to one if the time is after November 2014 and 0 otherwise. $X_{it}$ denotes a $k$-dimensional vector of control variables.
3 Empirical results
Table 1 lists the descriptive statistics for main variables used in the regression model, including the mean value, max value, min value, standard deviation and number of observations, respectively. We can see that the mean value of the premium is 68.3% which is consistent with previous research.
Table 1: Descriptive Statistics
Variable Mean Max Min Median SD N
Prem 0.683 7.759 -0.346 0.462 0.832 6,462
SHHKPolicy 0.558 1.000 0.000 1.000 0.497 6,462
Effboth -0.015 -0.002 -0.067 -0.014 0.006 6,462
Turnover 2.009 120.209 0.012 1.035 3.429 6,462
SOE 0.919 1.000 0.000 1.000 0.273 6,462
Demand 2.928 48.046 0.040 2.614 1.951 6,462
Size 25.028 28.555 21.262 24.906 1.456 6,462
Interest rate 9.020 23.380 1.355 9.095 5.414 6,462
3.1 Baseline Effect of SHHK Stock Connect
We first examines the effect of the SHHK Stock Connect on the A-H share price premium. Table 2 reports the baseline dynamic panel regression results. Column (1) only includes the lagged premium and the SHHK Stock Connect indicator. The coefficient on SHHKPolicy is positive and statistically significant at the 1% level, indicating that the introduction of the Stock Connect is associated with an increase in the A-H share price premium.
Columns (2) to (6) progressively incorporate additional firm-level control variables, with Column (6) representing the full specification. Across all specifications, the estimated coefficient on SHHKPolicy remains positive and significant at 1% level. In the full model, the coefficient on SHHKPolicy is 0.184, implying that, holding other variables constant, the implementation of the SHHK Stock Connect is associated with an average increase of approximately 18.4% in the A-H share price premium for a given pair of dual-listed shares. The baseline regression results present an average effect of the SHHK Stock Connect on the A-H share price premium. However, the economic mechanisms under this effect remain unclear. If SHHK Stock Connect operates by alleviating market frictions
Table 2: Baseline Effect of the SHHK Stock Connect on the A-H Share Price Premium
AH Premium AH Premium AH Premium AH Premium AH Premium AH Premium
(1) (2) (3) (4) (5) (6)
Lagged Premium 0.727∗∗∗ 0.676∗∗∗ 0.677∗∗∗ 0.676∗∗∗ 0.693∗∗∗ 0.694∗∗∗
(0.130) (0.144) (0.142) (0.142) (0.136) (0.136)
SHHKPolicy 0.184∗∗∗ 0.145∗∗∗ 0.144∗∗∗ 0.147∗∗∗ 0.180∗∗∗ 0.184∗∗∗
(0.017) (0.025) (0.024) (0.023) (0.022) (0.022)
Turnover 0.025 0.025 0.024 0.017∗∗ 0.017∗∗
(0.015) (0.015) (0.014) (0.009) (0.009)
SOE 0.055 0.053 0.030 0.030
(0.059) (0.062) (0.050) (0.049)
Demand -0.012 -0.005 -0.005
(0.014) (0.012) (0.012)
Size -0.075 -0.075
(0.044) (0.043)
Interest rate 0.002
Observations $63$ $63$ $63$ $63$ $63$ $63$
AR(2) $p$-value 0.315 0.912 0.915 0.891 0.797 0.758
Hansen Test $p$-value 0.548 0.416 0.417 0.425 0.462 0.469
  • Notes: This table reports dynamic panel estimates of the effect of the SHHK Stock Connect on the A-H share price premium. Columns (1)–(6) sequentially add firm-level control variables to the baseline specification. All models are estimated using two-step system GMM. The lagged dependent variable is instrumented using its own lags dated t2t-2 and t3t-3 in the differenced equation. Firm and year fixed effects are included. Robust two-step standard errors are reported in parentheses.
  • ∗∗∗, ∗∗, and denote statistical significance at the 1%, 5%, and 10% levels, respectively.
and improving information transmission, its impact should be more pronounced in less efficient markets. This motivates our analysis of efficiency-dependent effects in the next subsection.
3.2 Market Efficiency-Dependent Effects of the SHHK Stock Connect
Market efficiency plays a central role in determining how quickly asset prices reflect available information (Fama,, 1970, 1991). In the context of cross-listed shares, cross-market frictions and information-based barriers can impede arbitrage and contribute to deviations from price parity across trading venues (Gagnon and Andrew Karolyi,, 2010). More generally, when limits to arbitrage, trading frictions, or institutional constraints are present, prices of economically equivalent securities may deviate from parity and fail to converge quickly (Shleifer and Vishny,, 1997; Gagnon and Andrew Karolyi,, 2010).
Meng et al., (2023) finds that the SHHK Stock Connect increases the A-share market efficiency of AH dual listing firms through stock price informational efficiency and corporate governance. Motivated by their research, we test whether the effect of the SHHK Stock Connect on the A-H share price premium depends on market efficiency. We construct the following model:
$\text{Prem}_{i,t}=\alpha_{i}+\beta_{1}\text{Prem}_{i,t-1}+\beta_{2}\text{SHHKPolicy}_{t}+\beta_{3}\text{Effboth}_{t}+\beta_{4}\text{SHHKPolicy}\times\text{Effboth}_{t}+\gamma^{\prime}X_{it}+\epsilon_{it},$
where the $\text{Prem}_{i,t}$ is the A-H share price premium for share $i$ at month $t$. Follow Domowitz et al., (1997) and Fan and Wang, (2017), we control the lagged term $\text{Prem}_{i,t-1}$ to reduce the effect of autocorrelation.
[…]
Table 3: Efficiency-Dependent Effects of the SHHK Stock Connect
          AH Premium
          Lagged Premium           0.633∗∗∗
          SHHKPolicy           -0.490∗∗
          Effboth           39.554∗∗
          SHHKPolicy $\times$ Effboth           -45.900∗∗∗
          Turnover           0.022∗∗∗
          SOE           0.021
          Demand           -0.009
          Firm Size           -0.108∗∗∗
          Interest Rate           0.007∗∗∗
          Observations           6,363
          AR(2) $p$-value           0.211
          Hansen Test $p$-value           0.780
  • Notes: This table reports dynamic panel estimates examining how market efficiency conditions the effect of the SHHK Stock Connect on the AH price premium. $Effboth$ measures overall market efficiency across the A- and H-share markets, with higher values (less negative) indicating greater efficiency. The interaction term allows the policy effect to vary with market efficiency. The model is estimated using two-step system GMM. Lagged dependent variables and potentially endogenous regressors are instrumented using their own lags dated t2t-2 and t3t-3. Firm and year fixed effects are included. Robust standard errors are reported in parentheses.
  • ∗∗∗, ∗∗, and denote statistical significance at the 1%, 5%, and 10% levels, respectively.
isolation. In this model, the marginal effect of SHHK Stock Connect is:
$\frac{\partial\text{Prem}}{\partial\text{SHHKPolicy}}=\beta_{2}+\beta_{4}\text{Effboth},$
so that, under our scaling where $\texttt{Effboth}<0$ and values closer to zero indicate higher efficiency, the estimated policy effect increases as Effboth becomes more negative (i.e., as efficiency declines). Consequently, the Stock Connect is associated with a larger change in the AH premium for firms operating under less efficient market conditions than for those with higher efficiency. Overall, the results in Table 3 support an efficiency-dependent policy effect both statistically and economically.
3.3 Robustness check
In this subsection, we employ several robustness checks to make our result more reliable. Firstly, we examine whether the initial policy announcement generated an immediate pricing response. Specifically, we use April 2014 as an alternative event date, motivated by Premier Li Keqiang’s statement on 10 April 2014 regarding the establishment of a Shanghai-Hong Kong stock market connectivity mechanism. [^7] On 10 April 2014, Premier Li Keqiang stated that China would actively create conditions to establish a Shanghai-Hong Kong stock exchanges connectivity mechanism to promote further two-way opening-up and the healthy development of the mainland and Hong Kong capital markets. Table 4 reports GMM estimates across specifications in Columns (1)-(6), where control variables are sequentially added and Column (6) presents the full model. Across all columns, the estimated coefficient on Announcement is close to zero and statistically insignificant, suggesting no economically meaningful change in the AH premium around the announcement date at the monthly frequency. A potential interpretation is that the announcement did not immediately relax the trading, settlement, or capital-account constraints that limit cross-market arbitrage; consequently, any convergence forces induced by the policy were unlikely to materialize until implementation.
Table 4: Announcement Effect of the SHHK Stock Connect on the A-H Share Price Premium
AH Premium AH Premium AH Premium AH Premium AH Premium AH Premium
(1) (2) (3) (4) (5) (6)
Lagged Premium 0.788∗∗∗ 0.725∗∗∗ 0.726∗∗∗ 0.726∗∗∗ 0.748∗∗∗ 0.748∗∗∗
(0.115) (0.125) (0.124) (0.126) (0.115) (0.115)
Announcement 0.001 -0.008 -0.008 -0.007 0.004 0.004
(0.019) (0.022) (0.022) (0.021) (0.015) (0.017)
Turnover 0.023 0.023 0.022 0.016∗∗ 0.016∗∗
(0.013) (0.013) (0.012) (0.007) (0.007)
SOE 0.048 0.046 0.027 0.027
(0.048) (0.051) (0.039) (0.039)
Demand -0.009 -0.004 -0.004
(0.012) (0.009) (0.009)
Size -0.060 -0.060
(0.036) (0.036)
Interest rate -0.000
Observations $63$ $63$ $63$ $63$ $63$ $63$
AR(2) $p$-value 0.152 0.621 0.627 0.596 0.449 0.452
Hansen Test $p$-value 0.522 0.404 0.405 0.413 0.442 0.441
  • Notes: This table reports dynamic panel estimates of the effect of the announcement of SHHK Stock Connect on the A-H share price premium. Columns (1)–(6) sequentially add firm-level control variables to the baseline specification. All models are estimated using two-step system GMM. The lagged dependent variable is instrumented using its own lags dated t2t-2 and t3t-3 in the difference equation. Firm and year fixed effects are included. Robust two-step standard errors are reported in parentheses.
  • ∗∗∗, ∗∗, and denote statistical significance at the 1%, 5%, and 10% levels, respectively.
Table 5: Efficiency-Dependent Effects of the announcement of SHHK Stock Connect
          AH Premium
          Lagged Premium           0.678∗∗∗
          Announcement           -0.217
          Effboth           13.023
          Announcement $\times$ Effboth           -14.471
          Turnover           0.021∗∗∗
          SOE           0.026
          Demand           -0.005
          Size           -0.087∗∗
          Interest rate           0.001
          Observations           6,363
          AR(2) $p$-value           0.298
          Hansen Test $p$-value           0.778
  • Notes: This table reports dynamic panel estimates examining how market efficiency conditions the effect of the announcement of SHHK Stock Connect on the AH price premium. Effboth measures overall market efficiency across the A- and H-share markets, with higher values (less negative) indicating greater efficiency. The interaction term allows the policy effect to vary with market efficiency. The model is estimated using two-step system GMM. Lagged dependent variables and potentially endogenous regressors are instrumented using their own lags dated t2t-2 and t3t-3. Firm and year fixed effects are included. Robust standard errors are reported in parentheses.
  • ∗∗∗, ∗∗, and denote statistical significance at the 1%, 5%, and 10% levels, respectively.
Table 5 examines whether the announcement of the SHHK Stock Connect has an efficiency-dependent effect on the A-H price premium. The coefficient on Announcement is negative but statistically insignificant (0.217-0.217), suggesting no significant change in the AH premium around the announcement date. More importantly, the interaction term Announcement $\times$ Effboth is also statistically insignificant (14.471-14.471), providing no evidence that the announcement effect varies systematically with market efficiency. Consistent with these results, the main effect of Effboth is not statistically different from zero in this specification. Overall, Table 5 indicates that the announcement itself and the interaction with market efficiency are unlikely to generate a measurable impact on the AH premium.
According to Meng et al., (2023), we use an alternative market efficiency proxy, Effboth2, as a robustness check. Specifically, for each trading day $t$ we compute the relative high-low spread as
$\text{S}_{2}=\frac{\text{H}_{t}-\text{L}_{t}}{\big(\text{H}_{t}-\text{L}_{t}\big)/2},$ (4)
and then take the month average of this daily measure. Similar to our previous definition regarding market efficiency, we use the negative spread to be a proxy for market efficiency. That is:
$\text{Eff}_{2}=-\text{S}_{2},$ (5)
For A-share market and H-share market, we calculate the $\texttt{Eff}_{2}$ separately and then add them up as an alternative proxy for overall market efficiency (EFFboth2). Table 6 presents our regression results for third robustness check. Even though SHHKPolicy is not statistically significant, the sign of coefficient is negative. The interaction term $\texttt{SHHKPolicy}\times\texttt{Effboth2}$ is negatively significant at 5% level. So, in general, this result is consistent with the result shown in Table 3. The SHHK Stock Connect was launched in November, 2014, to further test whether the effect comes from SHHK Stock Connect and assess the robustness of our previous findings, we conduct a placebo test by assigning the implementation of the SHHK Stock Connect to a later date. Specifically, we assume that the policy was introduced two years after its actual launch and construct a binary indicator, FakePolicy, which equals one
Table 6: Efficiency-Dependent Effects of the SHHK Stock Connect using alternative efficiency measurement
          AH Premium
          Lagged Premium           0.796∗∗∗
          SHHKPolicy           -0.130
          Effboth2           2.370
          SHHKPolicy $\times$ Effboth2           -4.504∗∗
          Turnover           0.011∗∗
          SOE           0.033
          Demand           0.000
          Size           -0.042∗∗
          Interest rate           0.006∗∗∗
          Observations           $63$
          AR(2) $p$-value           0.406
          Hansen Test $p$-value           0.272
  • Notes: This table reports dynamic panel estimates examining how market efficiency conditions the effect of the SHHK Stock Connect on the AH price premium. Effboth2 measures overall market efficiency across the A- and H-share markets, with higher values (less negative) indicating greater efficiency. The interaction term allows the policy effect to vary with market efficiency. The model is estimated using two-step system GMM. Lagged dependent variables and potentially endogenous regressors are instrumented using their own lags dated t2t-2 and t3t-3. Firm and year fixed effects are included. Robust standard errors are reported in parentheses.
  • ∗∗∗, ∗∗, and denote statistical significance at the 1%, 5%, and 10% levels, respectively.
for observations after November 2016 and zero otherwise.
Table 7 reports the placebo regression results for the effect of SHHK Stock Connect on share price premium. In contrast to the baseline estimates, the coefficient on FakePolicy is positive but insignificant across all columns. This finding suggests that share price premium is not driven by common time trend, the SHHK stock policy does affect the premium. Next step, we conducted the placebo test regarding whether there is an efficiency dependent effects of the SHHK Stock Connect. Table 8 reports a placebo test in which the SHHK Stock Connect is assumed to have been implemented two years later, in November 2016. The key coefficient of interest is the interaction term between the placebo policy indicator and market efficiency, $\texttt{FakePolicy}\times\texttt{Effboth}$.
The interaction term $\texttt{FakePolicy}\times\texttt{Effboth}$ is statistically insignificant, indicating that when the policy timing is artificially shifted, market efficiency does not systematically condition the relationship between the placebo policy and the AH price premium. This result suggests that the efficiency-dependent effect documented in the previous analysis is not driven by common time trends or spurious correlations, but is instead specific to the actual implementation of the SHHK Stock Connect.
Table 7: Placebo test
AH Premium AH Premium AH Premium AH Premium AH Premium AH Premium
(1) (2) (3) (4) (5) (6)
Lagged Premium 0.790∗∗∗ 0.725∗∗∗ 0.726∗∗∗ 0.727∗∗∗ 0.751∗∗∗ 0.751∗∗∗
(0.118) (0.126) (0.125) (0.128) (0.117) (0.118)
FakePolicy 0.013 0.000 0.001 0.001 0.011 0.011
(0.030) (0.036) (0.036) (0.036) (0.031) (0.032)
Turnover 0.023 0.023 0.022 0.016∗∗ 0.016∗∗
(0.013) (0.013) (0.012) (0.007) (0.007)
SOE 0.049 0.047 0.027 0.027
(0.047) (0.050) (0.039) (0.038)
Demand -0.009 -0.004 -0.004
(0.011) (0.009) (0.009)
Size -0.059 -0.059
(0.036) (0.036)
Interest rate -0.000
Observations $63$ $63$ $63$ $63$ $63$ $63$
AR(2) $p$-value 0.158 0.615 0.621 0.589 0.450 0.452
Hansen Test $p$-value 0.527 0.402 0.403 0.412 0.445 0.445
  • Notes: This table reports placebo tests in which the SHHK Stock Connect is assumed to have been implemented two years later, in November 2016. FakePolicy is a binary indicator equal to one for observations after the placebo date and zero otherwise. The dependent variable is the monthly AH price premium. The model is estimated using two-step system GMM. Lagged dependent variables are instrumented using their own lags dated t2t-2 and t3t-3. Firm and year fixed effects are included. Robust standard errors are reported in parentheses.
  • ∗∗∗, ∗∗, and denote statistical significance at the 1%, 5%, and 10% levels, respectively.
Table 8: Efficiency-Dependent Effects of the SHHK Stock Connect using alternative efficiency measurement
          AH Premium
          Lagged Premium           0.680∗∗∗
          FakePolicy           0.150
          Effboth           -1.852
          FakePolicy $\times$ Effboth           11.928
          Turnover           0.019∗∗∗
          SOE           0.026
          Demand           -0.007
          Size           -0.080∗∗
          Interest rate           -0.001
          Observations           $63$
          AR(2) $p$-value           0.466
          Hansen Test $p$-value           0.849
  • Notes: This table reports placebo tests in which the SHHK Stock Connect is assumed to have been implemented two years later, in November 2016. FakePolicy is a binary indicator equal to one for observations after the placebo date and zero otherwise. The dependent variable is the monthly AH price premium. The model is estimated using two-step system GMM. Lagged dependent variables are instrumented using their own lags dated t2t-2 and t3t-3. Firm and year fixed effects are included. Robust standard errors are reported in parentheses.
  • ∗∗∗, ∗∗, and denote statistical significance at the 1%, 5%, and 10% levels, respectively.
4 Conclusion and Limitation
This study examines how the Shanghai–Hong Kong Stock Connect affects the A–H share price premium and whether the policy impact varies with market efficiency. Our findings are as follows: Firstly, the implementation of the Shanghai–Hong Kong Stock Connect significantly increases the A-H premium by 18.4% on average. Second, the effect is heterogeneous and depends on the overall market efficiency. When we interact the policy indicator with market-efficiency measure, the marginal policy effect becomes pronounced for firms operating under lower market efficiency conditions, which increases the A-H premium; on the other hand, the SHHK Stock Connect can reduce the A-H premium for the dual-listing firms operating in higher market efficiency. Finally, we find no evidence of a premium response at the announcement stage. When April 2014 is used as an announcement event date, the announcement indicator and its interaction with efficiency are statistically insignificant, suggesting that the A-H premium is unlikely to change prior to the implementation of the SHHK Stock Connect. This study also has two limitations. First, our sample is restricted to A–H firms whose A-shares are listed in Shanghai, so the evidence should be interpreted as specific to SHHK-eligible firms and may not generalise to Shenzhen-listed dual-listed firms. So, in the future, we may expand the sample to incorporate Shenzhen-listed A–H firms and examine whether the Shenzhen–Hong Kong Stock Connect exhibits a similar pattern. Second, while we find the Shanghai Hong Kong Stock Connect has different effects across firms, the mechanism explaining why the premium increases remains underdeveloped. Future research could introduce more direct channel variables, such as northbound and southbound holdings shares.